OpenStudy (anonymous):

csc^3(-x) + 8 divided by csc(x) - 2 becomes 8-csc^3(x) divided by -2-csc(x). I don't get why -2-csc(x)? it should be -2 + csc(x) if you just rearrange it like the numerator but I see that this way you cannot cancel across.

OpenStudy (anonymous):

what do you want to cancel off?

OpenStudy (anonymous):

the final answer is (2-cscx) (4+2cscx + csc^2x) divided by -(2-cscx). Both 2-cscx cancel out so you are left with -(4+2cscx + csc^2x)

OpenStudy (anonymous):

difference of two cubes was used

OpenStudy (anonymous):

oh.. csc(-x) = -csc(x) and cube of a negative is negative

OpenStudy (anonymous):

your numerator is in fact, -csc^3(x)+2

OpenStudy (anonymous):

-csc^3(x)+2^3

OpenStudy (anonymous):

i get that but csc (x) is in the denominator so csc (x) -2 how does it become -2-cscx??

OpenStudy (anonymous):

\[ \frac{-\csc^3(x)+8}{\csc(x)-2}=-\frac{(\csc(x)-2)(\csc^2(x)+2\csc(x)+4)}{\csc(x)-2}=-[\csc^2(x)+2\csc(x)+4] \]

OpenStudy (anonymous):

@electrokid csc^3 (-x). not, -csc^3 x

OpenStudy (anonymous):

thats what you said is the answer!

OpenStudy (anonymous):

yes, you are right eledtrokid.

OpenStudy (anonymous):

i know. :)

OpenStudy (anonymous):

yes.... \[\csc^3(−x)=[\csc(−x)]^3=[−\csc(x)]^3=−\csc^3(x)\]

OpenStudy (anonymous):

hihihi, but the answer is trivial , right? just plug your logic in both numerator and denominator, and both sides are equal., right?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

the numerator = 8-xsc^3 just rearrange the order.

OpenStudy (anonymous):

the same with denominator. so get the answer. am I right?

OpenStudy (anonymous):

hey guys, where are both you?

OpenStudy (anonymous):

oh yea..

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

hihihi... that 's the " desease " of the high level students, thinking too far from the ground

OpenStudy (anonymous):

there you go.. we should learn to un-learn things..

OpenStudy (anonymous):

don't understand

OpenStudy (anonymous):

see what I mean.. you are already half way up there.. lol

OpenStudy (anonymous):

@Hoa we work so hard to learn things. It is much harder to "un-learn" things and is never taught.

OpenStudy (anonymous):

got what you mean, hihihi... since I don't have a giant brain to think big, just think as small as I am. by the way, you seem very good at trigs? right? let me become your fan to get benefit in future

OpenStudy (anonymous):

@Hoa If you get a chance, get hold of the book "Escher, Godel and Bach: The eternal golden Braid" by Douglas Hofstader. You'll see exactly what I mean.

OpenStudy (anonymous):

I will. but not now. too many things to do with my upcoming test on next Thursday

OpenStudy (anonymous):

ohoh, you introduce me a wordplaying game book. how can i understand with my bad English? I am immigrant. not native, but i will try to readd. thanks anyway